OpenStudy (king):

A circle and square have same perimeter.Then: a)areas are equal b)area of circle is greater c)area of square is greater d)area of circle is pi times greater than area of square Please explain

OpenStudy (king):

???????????????????????????????????????????????

OpenStudy (anonymous):

\[A_{square} = (\frac{P}{4})^2\] \[A_{circle} = \pi*(\frac{P}{2\pi})^2\] looks like the area of the circle is greater

OpenStudy (king):

yeah but by how much

OpenStudy (king):

which option and why

OpenStudy (anonymous):

option B. seems the area of a circle is \[\frac{4}{\pi}\]times greater than the area of a square

OpenStudy (king):

no u chose option wrong but ure method is right it will be\[P ^{2}\div4\pi ^{2}\times \pi so that will be P/4\pi\] so option d

OpenStudy (king):

ok thnx agd see the other question

OpenStudy (anonymous):

how could my option be wrong?

OpenStudy (anonymous):

agdgdgdgwngo is right